# WZW terms without anomalies: generalised symmetries in chiral Lagrangians

### v1

@@ 5.1 Non-minimal scalar charge
But what if the scalar charge is taken to be non-minimal, $|X_{\phi}|\neq 1$? This naïvely presents a puzzle, because the 2-group structure is not modified (depending only on chiral fermion representations in the UV theory), but the coefficient of the WZW does appear to be rescaled by $X_{\phi}$ upon matching, since the vertex coupling the gauge field to the scalars is rescaled.
This requires a slightly more careful analysis, because the non-minimal scalar charge alters the symmetry breaking structure induced by the scalars condensing. Importantly, if the scalars have charge $|X_{\phi}|\neq 1$ under the gauge field (for which we now take the quarks to have charge $+1$), the $\mathrm{U}(1)$ gauge symmetry is broken only down to a discrete $\mathbb{Z}_{|X_{\phi}|}\subset\mathrm{U}(1)$ by the condensate. The target space arising from the global symmetry breaking in turn becomes the quotient
$S^{2}/\mathbb{Z}_{|X_{\phi}|}\,,$ (76)
for now ignoring the QCD part which is unchanged. This manifold is diffeomorphic to another 2-sphere (albeit with singular points appearing in the quotient metric at the poles), but the volume form must be rescaled,
$\rm{Vol}_{S^{2}/\mathbb{Z}_{|X_{\phi}|}}=X_{\phi}\,\mathrm{Vol}_{S^{2}}$ (77)
in order to remain integral (since we effectively integrate only over a wedge defined by opening angle $\Delta\phi=2\pi/|X_{\phi}|$ to obtain the volume of the ‘quotiented’ sphere). The object $j_{\phi}$ obtained from the Lagrangian is clearly rescaled as $j_{\phi}\mapsto X_{\phi}j_{\phi}$, and so this factor drops out when we trade the curvature $dj_{\phi}$ for the volume form, that is $dj_{\phi}=-4\pi v^{2}\mathrm{Vol}_{S^{2}}$ remains true.
The upshot is that the coefficient of the mixed WZW term does not depend on $X_{\phi}$ (not even on its sign). This is consistent with matching the 2-group symmetry, resolving our little puzzle.
@@ 5.3 From $\mathbb{C}P^{1}$ to $\mathbb{C}P^{n}$
$H^{\bullet}_{\mathrm{dR}}(\mathbb{C}P^{n})=\begin{cases}\mathbb{R}\qquad&n\text{~{}even~{}},0\leq n\leq n\\ 0\qquad&n\text{~{}odd~{}}\,,\end{cases}$ (81)
@@
Yet further generalisation is possible if we allow $K$ to be a more general manifold $M$, not necessarily a coset arising from spontaneous symmetry breaking. The same mixed WZW term exists whenever $H^{2}_{\mathrm{dR}}(M)$ is non-trivial, with $\Omega$ similarly picked to be a 2-form representative of a non-trivial element in $H^{2}_{\mathrm{dR}}(M)$ with unit period. A partial UV-completion is provided by QCD coupled to a non-linear sigma model with target space being the line bundle $L\to M$ over $M$ whose Chern class is given by $\Omega$. We then gauge the diagonal between the $\mathrm{U}(1)_{q}$ quark number symmetry and the $\mathrm{U}(1)$ symmetry acting on the fibre of $L$ [^17] We thank Y. Tachikawa for suggesting this more general approach. The crucial non-trivial 2-group structure then arises from the mixed ’t Hooft anomaly between $\mathrm{U}(1)_{q}$ and $\text{SU}(N)_{L/R}$ after this gauging Cordova:2018cvg (the same phenomenon also happens when gauging finite groups with mixed ’t Hooft anomalies Tachikawa:2017gyf ).
@@ 6.2 Non-invertible symmetry from the UV anomaly polynomial
with $\beta$ being rational. [^19] For a slightly different take on the range of the transformation’s parameter, see Refs. GarciaEtxebarria:2022jky ; Karasik:2022kkq . The ill-defined CS terms are shorthands for 3d topological quantum field theories (TQFTs) localised on the defect submanifold $M_{3}$ which couple to the $\mathrm{U}(1)$ fields. For example, if we take $\beta=1/6n_{c}K$, the symmetry defect becomes

### v2

@@ 5.1 Non-minimal scalar charge
But what if the scalar charge is taken to be non-minimal, $X_{\phi}\neq 1$ (in minimal units, i.e. such that $\text{gcd}(X_{q},X_{\phi})=1$)? This naïvely presents a puzzle, because the 2-group structure is not modified, depending only on chiral fermion representations in the UV theory, but the coefficient of the WZW does appear to be rescaled, this time by a factor $1/X_{\phi}$ that arises due to the rescaling of both the $b\phi_{i}\phi_{i}^{\dagger}$ vertex and of the gauge boson mass. This scenario requires a slightly more careful analysis, because the non-minimal scalar charge alters the symmetry breaking structure induced by the scalars condensing.
Firstly, the non-minimal scalar charge means there is a discrete $\mathbb{Z}_{|X_{\phi}|}\subset\mathrm{U}(1)$ gauge symmetry that remains unbroken by the scalar condensate, which acts non-trivially on the quark fields and can in principle be detected via its holonomy.
Concerning the dark pions, the target space remains a 2-sphere, but with a rescaled volume form, satisfying
$\int_{S^{2}}\mathrm{Vol}_{S^{2}}=X_{\phi}\,.$ (80)
To see this, recall that $\mathrm{Vol}_{S^{2}}=d\mathcal{A}_{\phi}$ where $\mathcal{A}_{\phi}=-i(dz_{i}^{*}z_{i}-z_{i}^{*}dz_{i})$ behaves like a gauge connection. However, under the $\mathrm{U}(1)$ gauge transformation $z_{i}\mapsto e^{iX_{\phi}\alpha}z_{i}$, with $\alpha$ being $2\pi$-periodic, we have,
$\mathcal{A}_{\phi}\mapsto\mathcal{A}_{\phi}-X_{\phi}d\alpha,$ (81)
instead of (68). Therefore, the properly normalised $\mathrm{U}(1)$ connection is $\mathcal{A}_{\phi}/X_{\phi}$ instead of $\mathcal{A}_{\phi}$. Consequently, the minimally quantised volume form is $\mathrm{Vol}_{S^{2}}/X_{\phi}$.
The WZW term can be derived in exactly the same way as in §4.3 by integrating out the massive $\mathrm{U}(1)$ gauge field, with a couple of changes already mentioned above. Firstly, since the mass squared of the gauge field is now equal to $X_{\phi}^{2}v^{2}$, the cross interaction term between $j_{B}$ and $j_{\phi}$ that arises from integrating out the gauge field is now inversely proportional to $X_{\phi}^{2}$:
$S_{\text{int}}=-\frac{n_{c}}{2X_{\phi}^{2}v^{2}}\int\star j_{B}\wedge j_{\phi}\,.$ (82)
Secondly, because $\phi$ now has charge $X_{\phi}$ under the $\mathrm{U}(1)$ gauge group, the current $j_{\phi}$ is rescaled:
$j_{\phi}=X_{\phi}j_{\phi}^{\text{(old)}}=2X_{\phi}v^{2}\mathcal{A}_{\phi}\,.$ (83)
Putting the two together, we arrive at
$S_{\text{int}}=2\pi n_{c}X_{q}\int_{\Sigma_{5}}\frac{1}{24\pi^{2}}\mathrm{Tr~{}}\left(g^{-1}dg\right)^{3}\wedge\frac{\mathrm{Vol}_{S^{2}}}{X_{\phi}}.$ (84)
Despite the modified coefficient, because of the compensating modified flux relation (80) the closed 5-form in the integrand is still integral, meaning this topological term is still of course well-defined. Moreover, the 2-group current algebra is unchanged i.e. is independent of $X_{\phi}$. To see this, we should take the Noether current for the 1-form symmetry to be $j^{(2)}_{\mathrm{wind}}:=\star_{4}\mathrm{Vol}_{S^{2}}/X_{\phi}$, which is integral, and so the minimal coupling term (14) is now $S_{\text{coup}}=i\int_{\Sigma_{4}}\frac{\mathrm{Vol}_{S^{2}}}{X_{\phi}}\wedge B$. If we repeat the analysis of §2, we see the Postnikov class of the 2-group remains $\hat{\kappa}_{L}=-\hat{\kappa}_{R}=n_{c}X_{q}$. This is consistent with matching the 2-group symmetry, resolving our little puzzle.
@@ 5.3 From $\mathbb{C}P^{1}$ to $\mathbb{C}P^{n}$
$H^{k}_{\mathrm{dR}}(\mathbb{C}P^{n})=\begin{cases}\mathbb{R}\qquad&k\text{~{}even~{}},0\leq k\leq n\\ 0\qquad&k\text{~{}odd~{}}\,,\end{cases}$ (88)
@@
Yet further generalisation is possible if we allow $K$ to be a more general manifold $M$, not necessarily a coset arising from spontaneous symmetry breaking. The same mixed WZW term exists whenever $H^{2}_{\mathrm{dR}}(M)$ is non-trivial, with $\Omega$ similarly picked to be a 2-form representative of a non-trivial element in $H^{2}_{\mathrm{dR}}(M)$ with unit period. A partial UV-completion is provided by QCD coupled to a non-linear sigma model with target space being the line bundle $L\to M$ over $M$ whose Chern class is given by $[\Omega]\in H^{2}(M;\mathbb{Z})$. We then gauge the diagonal between the $\mathrm{U}(1)_{q}$ quark number symmetry and the $\mathrm{U}(1)$ symmetry acting on the fibre of $L$ [^17] We thank Y. Tachikawa for suggesting this more general approach. The crucial non-trivial 2-group structure then arises from the mixed ’t Hooft anomaly between $\mathrm{U}(1)_{q}$ and $\text{SU}(N)_{L/R}$ after this gauging Cordova:2018cvg (the same phenomenon also happens when gauging finite groups with mixed ’t Hooft anomalies Tachikawa:2017gyf ).
@@ 6.2 Non-invertible symmetry from the UV anomaly polynomial
with $\beta$ being rational. [^19] For a slightly different take on the range of the transformation’s parameter, see Refs. Karasik:2022kkq ; GarciaEtxebarria:2022jky ; Arbalestrier:2024oqg . The ill-defined CS terms are shorthands for 3d topological quantum field theories (TQFTs) localised on the defect submanifold $M_{3}$ which couple to the $\mathrm{U}(1)$ fields. For example, if we take $\beta=1/6n_{c}K$, the symmetry defect becomes